OpenStudy (anonymous):

solve using the addition method 6x+3y=0 8x+5y=8

OpenStudy (anonymous):

You just add each together. 6x+8x=14x, 3y+5y=8y, 0+8=8 So you end up with 14x+8y=8 Did you need it in slope-intercept form?

OpenStudy (anonymous):

If you do, you subtract the x value over to get 8y=-14x+8 Then divide both sides by 8 to isolate y. y=-14/8x+1 Simplified it is y=-7/4x+1